# LLM-Enhanced Bayesian Optimization for Efficient Analog Layout Constraint Generation

### v1

@@ I Introduction
Recent progress in analog circuit dimensioning has significantly improved the effectiveness of leveraging simulations for performance enhancement [3, 6, 7, 8, 9]. These simulation-driven methods approach analog dimensioning as an optimization problem that does not reveal its internal workings, relying on circuit simulators to evaluate performance metrics. This approach requires minimal preliminary knowledge about circuit configurations, contrasting with model-dependent techniques that necessitate a deep understanding of complex design and performance interactions. The integration of Gaussian Process (GP)-based Bayesian Optimization (BO) for automatic transistor sizing [10, 11] combines the strengths of both paradigms, achieving significant reductions in simulation requirements while maintaining adaptability across various circuit designs. This advancement highlights the importance of incorporating iterative simulations into the design feedback loop, inspiring the development of an integrated, feedback-oriented analog layout synthesizer.
@@
    All the algorithms, experiments and benchmarks are open-sourced in the LLANA framework, publicly available at https://anonymous.4open.science/r/LLANA-2F4B.
@@ III-B Large Language Models (LLMs)-enhanced BO
BO’s effectiveness depends on the quality of surrogate models and sampling strategies, challenged by limited data, sensitivity to inaccuracies, and the difficulty of incorporating prior knowledge into new tasks. Recently, there has been a growing trend of utilizing AI and LLMs to address challenges within the EDA domain [15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41]. [42] explained LLM in-context learning (ICL) as performing implicit Bayesian inference [12]. LLMs can enhance BO by leveraging: (1) prior knowledge through ICL for tapping into pre-trained insights [43], (2) the ability to generalize from limited examples, aiding in efficient exploration [44], and (3) processing contextual information to enrich optimization tasks and search strategies [45].

### v3

@@ I Introduction
Recent progress in analog circuit dimensioning has significantly improved the effectiveness of leveraging simulations for performance enhancement [3, 6, 7, 8, 9]. These simulation-driven methods approach analog dimensioning as an optimization problem that does not reveal its internal workings, relying on circuit simulators to evaluate performance metrics. This approach requires minimal preliminary knowledge about circuit configurations, contrasting with model-dependent techniques that necessitate a deep understanding of complex design and performance interactions. The integration of Gaussian Process (GP)-based Bayesian Optimization (BO) for automatic transistor sizing combines the strengths of both paradigms, achieving significant reductions in simulation requirements while maintaining adaptability across various circuit designs. This advancement highlights the importance of incorporating iterative simulations into the design feedback loop, inspiring the development of an integrated, feedback-oriented analog layout synthesizer.
@@
    All the algorithms, experiments and benchmarks are open-sourced at https://github.com/dekura/LLANA.
@@ III-B Large Language Models (LLMs)-enhanced BO
BO’s effectiveness depends on the quality of surrogate models and sampling strategies, challenged by limited data, sensitivity to inaccuracies, and the difficulty of incorporating prior knowledge into new tasks. Recently, there has been a growing trend of utilizing AI and LLMs to address challenges within the EDA domain [13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42]. [43] explained LLM in-context learning (ICL) as performing implicit Bayesian inference [10]. LLMs can enhance BO by leveraging: (1) prior knowledge through ICL for tapping into pre-trained insights [44], (2) the ability to generalize from limited examples, aiding in efficient exploration [45], and (3) processing contextual information to enrich optimization tasks and search strategies [46].